# Automated Compilation Including Dropouts: Tolerating Defective Components in Stabiliser Codes

### v1

@@ 5 Bivariate bicycle codes
Figure 8: Connectivities for the bivariate bicycle (BB) code. Vertices are qubits, and solid lines between them represent the presence of couplers. I’ve used the standard polynomial specification of BB codes where qubits are labelled as $q_{L/R}$, with $q$ a monomial in a finite polynomial ring and a left (‘L’) or ‘right’ (‘R’) label. a. The local connectivities of $X$ (blue) and $Z$ (pink/purple) stabilisers in construction taken from [shawLoweringConnectivityRequirements2025]. This construction results in 5-regular global connectivity graph. b. The local connectivity of the same stabilisers in my hexagonal connectivity. Note that the qubits associated with monomials $q_{L}$ and $q_{R}$ are no longer connected, but we introduce two new edges with directions $a_{2}b_{2}^{-1}$ and $a_{3}b_{3}^{-1}$. c. All of the stabilisers containing a given $L$ qubit in the hexagonal connectivity, showing the 6-regularity of the graph.
Notwithstanding discussions of possible theoretical methods for doing so [xuFaulttolerantProtocolsSpacetime2025, zhouLouvreRelaxingHardware2025], I believe ACID is the first concrete proposal for handling dropouts in the bivariate bicycle (BB) codes [bravyiHighthresholdLowoverheadFaulttolerant2024b]. Following [shawLoweringConnectivityRequirements2025], I applied ACID to the 144-qubit and 288-qubit codes, each with 12 logical qubits, from [bravyiHighthresholdLowoverheadFaulttolerant2024b] as mid-cycle codes. In the dropout-free case, these mid-cycle codes produce end-cycle codes with distance 6 and distance 12 respectively (these are the distances I use as the length of the memory experiments).
I applied ACID to these codes with two underlying connectivities, inspired by the hex-grid and square-grid surface code connectivities: first, the degree-5 connectivity introduced in [shawLoweringConnectivityRequirements2025], and second a ‘hexagonal’ connectivity in which the six qubits defining a stabiliser are connected in a cycle (see Figure 8). This latter connectivity is degree-6, but offers more flexibility in contraction schedules for each stabiliser. In particular, where in the degree-5 connectivity the loss of any coupler results in anticommuting quasi-stabilisers (as the local connectivity graph is a tree), the cyclic nature of the hexagonal connectivity means we can tolerate the loss of isolated couplers whilst still being able to measure all the original stabilisers. It is interesting to note that for some BB codes, such as the 288 code, the hexagonal connectivity is completely symmetrical in its couplers and qubits (the associated undirected graph is arc-transitive), whereas in others such as the 144 code the edges split into two or more edge orbits.
The non-locality and complexity of these connectivities mean that neither will likely prove a viable layout for practical quantum hardware. Future work should examine the dropout tolerance of proposals such as [eberhardtPruningQLDPCCodes2024, liangPlanarQuantumLowDensity2025] that aim for more hardware-friendly QLDPC codes.
@@
Of particular interest is the single-dropped-coupler case for the 144-qubit code with the hexagonal connectivity (Appendix D). I note that the introduction of a single defective coupler results in most cases in a substantially lower logical error rate than in the dropout-free case. Investigating further, one can calculate the distance of the end-cycle codes resulting from a given schedule and see that certain contraction schedules lead to an end-cycle-distance of 7 or even 8. Alternative contraction schedules with higher distances are also discussed (in the degree-5, dropout-free case) in the appendix of [shawLoweringConnectivityRequirements2025]. It is noteworthy that the dropout-free schedules are tightly bunched in logical error rate, suggesting that the heuristic metrics that the solver optimises for in fact push it into a schedule that is sub-optimal, and that the frustration introduced by a new constraint allows the solver to explore better options.
Whereas for the surface code lattice surgery is straightforwardly generalised to the case of patches with dropouts, applying ACID to the BB codes corrodes the symmetry of the code that is used in proposals for logical processing with the code. In particular, the code is no longer translationally invariant, so we appear to lose access to a large class of automorphisms used in [yoderTourGrossModular2025] to perform universal computation.
@@ 6 The colour code
I applied ACID to three flavours of colour code, with all data shown in Appendix D. The first, the degree-3 ancilla-free colour code is taken directly from [gidneyNewCircuitsOpen2023a] (where it is called the ‘middle-out’ circuit). This connectivity is naturally implemented on a hex-grid qubit layout. The local connectivity graphs in this connectivity are then hexagons, so we expect to often be able to measure the original stabilisers in the presence of dropped couplers. However, the tightly packed and overlapping nature of the colour code stabilisers means that the global scheduling problem is very sensitive to local restrictions, and a single dropped coupler is normally enough to force the scheduler to require four contraction layers.
@@
Finally, I tried a version of the colour code based on the ‘superdense’ syndrome extraction circuit also from [gidneyNewCircuitsOpen2023a]. Eight qubits are assigned to each stabiliser of the colour code, six outer qubits and two inner ones (these inner qubits were previously measurement ancillae). I then give ACID the task of scheduling four stabilisers per stabiliser of the colour code: an $X$ and a $Z$ weight-8 stabiliser and $X$ and $Z$ weight-2 stabilisers supported on the inner qubits. The natural solution found by the solver is then almost exactly the superdense syndrome extraction circuit from the literature, but whereas the original superdense circuit has three ‘incoming’ CNOTs preceding three ‘outgoing’ CNOTs, in this circuit the two groups take turns to go first. This leads to the same number of detectors, but they are more localised in time, suggesting the decoding performance may be better. This superdense colour code is embedded on a full square-grid lattice. This means we have two redundant edges in the local connectivity graphs of the weight-8 stabilisers. This redundancy, combined with the fact that the contraction circuits take four timesteps, can lead to a combinatorial explosion in the number of contraction schedules, overwhelming the solver (even though in principle having lots of schedules to choose from is a good thing). To overcome this, we prune the contraction schedules ahead of feeding them to the solver based on their compatibility with certain ‘preferred schedules’.
@@ 7 Outlook
Whilst the basic ‘merge and split’ operations of surface code and colour code lattice surgery [litinskiGameSurfaceCodes2019f, landahlQuantumComputingColorcode2014] should still be valid in the presence of dropouts, it is not clear how the loss of symmetry of codes in the presence of dropouts will affect logical operations for more general codes.
@@ Appendix A Distance definitions
where $L$ acts only on the logical qubits, and $G$ acts only on the gauge qubits, and we require that $L$ is non-trivial. We can calculate the ‘dressed distance’ of the subsystem code with $g$ gauge qubits by gauge-fixing. For each of the $2^{g}$ binary vectors $z$ of length $g$, we construct a stabiliser code with an $X$ parity check matrix given by the parity check matrix of the code $H^{X}$ with the rows of $G^{X}$ corresponding to the 1s in $z$ appended, and a $Z$ parity check matrix constructed by appending the rows of $G^{Z}$ where $z$ is 0 to $H^{Z}$. We find the distance of each of these codes using the QDistRand package [pryadkoQDistRndGAPPackage2022], and take the minimum.
@@ Appendix C Memory simulations
A quantum memory experiment should, as far as possible, simulate a portion of what would, in an actual quantum computation, be a period of qubit idling preceded and followed by either more idling or logical operations. The Stim [gidneyStimFastStabilizer2021b] memory experiments I ran to produce the plots in Appendix D have the following stages. For the noisy portion of the circuit, I assume $p=0.001$ probability of depolarisation after one- and two-qubit gates, and include bit- and phase-flips in the appropriate basis after resets and before measurements also with probability $p$.
@@ Appendix D Performance plots
All memory experiments were simulated at $p=0.001$ physical error rate and decoded with BP-AC [wolanskiAmbiguityClusteringAccurate2025a].
@@ Appendix E Reproducing LUCI
To compare ACID against LUCI [debroyLUCISurfaceCode2024, higgottHandlingFabricationDefects2025] in the surface code, I implemented a version of the LUCI algorithm as a special case of ACID. The product stabilisers, gauge and logical operators are found by the usual ACID routine, as are the local contraction schedules and the constraints between them.
Rather than tasking the solver with minimising the number of contraction layers required for a syndrome extraction circuit, we divide the surface code quasi-stabilisers into four ‘colours’ 1-4 as described in the original LUCI paper, and require that the solver contract and measure the quasi-stabilisers of a given colour $i$ on layer $i$. The solver is not required to use any particular schedule in measuring the required quasi-stabilisers, and is free to ‘fill in’ additional quasi-stabiliser contractions on each layer as far as possible. It is intended that this represents an upper bound on the flexibility of the LUCI algorithm. This implementation does not explicitly attempt to construct ‘shells’ as described in [higgottHandlingFabricationDefects2025]. However, it does attempt to measure quasi-stabilisers that form part of product stabilisers as much as possible, including ‘redundant’ contractions that are not strictly required to measure product stabilisers, and it automatically assigns additional detectors in these cases, which likely amounts to the same thing.

### v2

@@ 5 Bivariate bicycle codes
Figure 8: Connectivities for the bivariate bicycle (BB) code. Vertices are qubits, and solid lines between them represent the presence of couplers. I’ve used the standard polynomial specification of BB codes where qubits are labelled as $q_{L/R}$, with $q$ a monomial in a finite polynomial ring and a left (‘L’) or ‘right’ (‘R’) label. a. The local connectivities of $X$ (blue) and $Z$ (pink/purple) stabilisers in construction taken from [24]. This construction results in 5-regular global connectivity graph. b. The local connectivity of the same stabilisers in my hexagonal connectivity. Note that the qubits associated with monomials $q_{L}$ and $q_{R}$ are no longer connected, but we introduce two new edges with directions $a_{2}b_{2}^{-1}$ and $a_{3}b_{3}^{-1}$. c. All of the stabilisers containing a given $L$ qubit in the hexagonal connectivity, showing the 6-regularity of the graph.
Notwithstanding discussions of possible theoretical methods for doing so [30, 32], I believe ACID is the first concrete proposal for handling dropouts in the bivariate bicycle (BB) codes [6]. Following [24], I applied ACID to the 144-qubit and 288-qubit codes, each with 12 logical qubits, from [6] as mid-cycle codes. In the dropout-free case, these mid-cycle codes produce end-cycle codes with distance 6 and distance 12 respectively (these are the distances I use as the length of the memory experiments).
I applied ACID to these codes with two underlying connectivities, inspired by the hex-grid and square-grid surface code connectivities: first, the degree-5 connectivity introduced in [24], and second a ‘hexagonal’ connectivity in which the six qubits defining a stabiliser are connected in a cycle (see Figure 8). This latter connectivity is degree-6, but offers more flexibility in contraction schedules for each stabiliser. In particular, where in the degree-5 connectivity the loss of any coupler results in anticommuting quasi-stabilisers (as the local connectivity graph is a tree), the cyclic nature of the hexagonal connectivity means we can tolerate the loss of isolated couplers whilst still being able to measure all the original stabilisers. It is interesting to note that for some BB codes, such as the 288 code, the hexagonal connectivity is completely symmetrical in its couplers and qubits (the associated undirected graph is arc-transitive), whereas in others such as the 144 code the edges split into two or more edge orbits.
The non-locality and complexity of these connectivities mean that neither will likely prove a viable layout for practical quantum hardware. Future work should examine the dropout tolerance of proposals such as [9, 16] that aim for more hardware-friendly QLDPC codes.
@@
Of particular interest is the single-dropped-coupler case for the 144-qubit code with the hexagonal connectivity (Appendix D). I note that the introduction of a single defective coupler results in most cases in a substantially lower logical error rate than in the dropout-free case. Investigating further, one can calculate the distance of the end-cycle codes resulting from a given schedule and see that certain contraction schedules lead to an end-cycle-distance of 7 or even 8. Alternative contraction schedules with higher distances are also discussed (in the degree-5, dropout-free case) in the appendix of [24]. It is noteworthy that the dropout-free schedules are tightly bunched in logical error rate, suggesting that the heuristic metrics that the solver optimises for in fact push it into a schedule that is sub-optimal, and that the frustration introduced by a new constraint allows the solver to explore better options.
Whereas for the surface code lattice surgery is straightforwardly generalised to the case of patches with dropouts, applying ACID to the BB codes corrodes the symmetry of the code that is used in proposals for logical processing with the code. In particular, the code is no longer translationally invariant, so we appear to lose access to a large class of automorphisms used in [31] to perform universal computation.
@@ 6 The colour code
I applied ACID to three flavours of colour code, with all data shown in Appendix D. The first, the degree-3 ancilla-free colour code is taken directly from [10] (where it is called the ‘middle-out’ circuit). This connectivity is naturally implemented on a hex-grid qubit layout. The local connectivity graphs in this connectivity are then hexagons, so we expect to often be able to measure the original stabilisers in the presence of dropped couplers. However, the tightly packed and overlapping nature of the colour code stabilisers means that the global scheduling problem is very sensitive to local restrictions, and a single dropped coupler is normally enough to force the scheduler to require four contraction layers.
@@
Finally, I tried a version of the colour code based on the ‘superdense’ syndrome extraction circuit also from [10]. Eight qubits are assigned to each stabiliser of the colour code, six outer qubits and two inner ones (these inner qubits were previously measurement ancillae). I then give ACID the task of scheduling four stabilisers per stabiliser of the colour code: an $X$ and a $Z$ weight-8 stabiliser and $X$ and $Z$ weight-2 stabilisers supported on the inner qubits. The natural solution found by the solver is then almost exactly the superdense syndrome extraction circuit from the literature, but whereas the original superdense circuit has three ‘incoming’ CNOTs preceding three ‘outgoing’ CNOTs, in this circuit the two groups take turns to go first. This leads to the same number of detectors, but they are more localised in time, suggesting the decoding performance may be better. This superdense colour code is embedded on a full square-grid lattice. This means we have two redundant edges in the local connectivity graphs of the weight-8 stabilisers. This redundancy, combined with the fact that the contraction circuits take four timesteps, can lead to a combinatorial explosion in the number of contraction schedules, overwhelming the solver (even though in principle having lots of schedules to choose from is a good thing). To overcome this, we prune the contraction schedules ahead of feeding them to the solver based on their compatibility with certain ‘preferred schedules’.
@@ 7 Outlook
Whilst the basic ‘merge and split’ operations of surface code and colour code lattice surgery [18, 14] should still be valid in the presence of dropouts, it is not clear how the loss of symmetry of codes in the presence of dropouts will affect logical operations for more general codes.
@@ Appendix A Distance definitions
where $L$ acts only on the logical qubits, and $G$ acts only on the gauge qubits, and we require that $L$ is non-trivial. We can calculate the ‘dressed distance’ of the subsystem code with $g$ gauge qubits by gauge-fixing. For each of the $2^{g}$ binary vectors $z$ of length $g$, we construct a stabiliser code with an $X$ parity check matrix given by the parity check matrix of the code $H^{X}$ with the rows of $G^{X}$ corresponding to the 1s in $z$ appended, and a $Z$ parity check matrix constructed by appending the rows of $G^{Z}$ where $z$ is 0 to $H^{Z}$. We find the distance of each of these codes using the QDistRand package [22], and take the minimum.
@@ Appendix C Memory simulations
A quantum memory experiment should, as far as possible, simulate a portion of what would, in an actual quantum computation, be a period of qubit idling preceded and followed by either more idling or logical operations. The Stim [11] memory experiments I ran to produce the plots in Appendix D have the following stages. For the noisy portion of the circuit, I assume $p=0.001$ probability of depolarisation after one- and two-qubit gates, and include bit- and phase-flips in the appropriate basis after resets and before measurements also with probability $p$.
@@ Appendix D Performance plots
All memory experiments were simulated at $p=0.001$ physical error rate and decoded with BP-AC [29].
@@ Appendix E Reproducing LUCI
To compare ACID against LUCI [8, 13] in the surface code, I implemented a version of the LUCI algorithm as a special case of ACID. The product stabilisers, gauge and logical operators are found by the usual ACID routine, as are the local contraction schedules and the constraints between them.
Rather than tasking the solver with minimising the number of contraction layers required for a syndrome extraction circuit, we divide the surface code quasi-stabilisers into four ‘colours’ 1-4 as described in the original LUCI paper, and require that the solver contract and measure the quasi-stabilisers of a given colour $i$ on layer $i$. The solver is not required to use any particular schedule in measuring the required quasi-stabilisers, and is free to ‘fill in’ additional quasi-stabiliser contractions on each layer as far as possible. It is intended that this represents an upper bound on the flexibility of the LUCI algorithm. This implementation does not explicitly attempt to construct ‘shells’ as described in [13]. However, it does attempt to measure quasi-stabilisers that form part of product stabilisers as much as possible, including ‘redundant’ contractions that are not strictly required to measure product stabilisers, and it automatically assigns additional detectors in these cases, which likely amounts to the same thing.